OpenStudy (anonymous):

Is my answer right???? Multiply: (6+√5)(4-2√5) Answer: 24-10√5

OpenStudy (microbot):

i will tell you if you show me every step you did to reach that answer:)

OpenStudy (anonymous):

I found the answer to this, thank you :)

OpenStudy (anonymous):

I have another one though, I'll show you how I got it :)

OpenStudy (microbot):

ok:) nice ill wait for it

OpenStudy (anonymous):

Thanks!!! :)

OpenStudy (anonymous):

Which of the following is a solution to √x + 10 + 8 = 3 My answer: -5 How? √-5 + 10 + 8 = 3.60555127546

OpenStudy (microbot):

what are the other options you have?

OpenStudy (anonymous):

x = 15 x = 5 x = -5 x = -15

OpenStudy (microbot):

why do you think this: How? √-5 + 10 + 8 = 3.60555127546 proves that -5 is the solution to √x + 10 + 8 = 3? since you have to find a solution for √x + 10 + 8 = 3 you have to solve it for x.

OpenStudy (microbot):

try to solve the equation for x

OpenStudy (anonymous):

I saw the three, and figured the rest would be cut off. Now i see how irrational it sounds.

OpenStudy (microbot):

15,-15,5 and -5 are all the options you have?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

OH1 It would be a positive, so that narrows it down

OpenStudy (microbot):

hmm well my answer would be i*sqrt(15) but there is no such choice....have they tought you the imaginary numbers yet?

OpenStudy (anonymous):

That's what I just got, the fifteen, but calculated there's a big number that doesn't even have the three in place

OpenStudy (anonymous):

it has been addressed briefly

OpenStudy (microbot):

no w8 i got it wrong ...give me a sec

OpenStudy (anonymous):

I did another problem just like it, and got one solid answer

OpenStudy (microbot):

did you get 225?

OpenStudy (microbot):

have to restart my modem brb!

OpenStudy (anonymous):

ok!

OpenStudy (microbot):

a question!: maybe the equation is not √x + 10 + 8 = 3 but √(x + 10) + 8 = 3? because like this it would make sence

OpenStudy (anonymous):

Let me check

OpenStudy (anonymous):

It's still the former

OpenStudy (anonymous):

√x + 10 = 3

OpenStudy (microbot):

without 8?

OpenStudy (anonymous):

OOPS!!!!!! wrong wrong wrong!

OpenStudy (anonymous):

√x + 10 = 8 = 3

OpenStudy (microbot):

write it again

OpenStudy (anonymous):

Honestly none of them seem to work

OpenStudy (anonymous):

Sorry i pressed the equal sign twice, my apologies

OpenStudy (anonymous):

√x + 10 + 8 = 3

OpenStudy (anonymous):

considering it equals a positive three, the only option are A and B, so I'm going with the former

OpenStudy (microbot):

i dont understand your way of deciding the answer...

OpenStudy (anonymous):

It makes sense by what I have on my paper, I'm sorry

OpenStudy (microbot):

you dont have to be sry...im just curious myself:)

OpenStudy (anonymous):

process of elimination, dear

OpenStudy (anonymous):

I see the three, that's better than getting an answer that's not even in the problem that GIVES you the answer (it being 3) also, it's a positive three Although it makes little to no sense, that's the way to go because there's either an error in the whole problem, or I'm reading too into it

OpenStudy (microbot):

i can tell you that if the equation were: √(x + 10) + 8 = 3 then: √(x + 10) = 3-8 √(x + 10) = -5 √(x + 10) =√((-5)^2) x+10=(-5)^2 x+10=25 x=25-10 x=15

OpenStudy (microbot):

anyway , when you know the right answer please let me know about it!:)GL with it!

OpenStudy (anonymous):

thanks!!!!!!! :) i will :) @MicroBot